how do I solve these algebraic equations : 1/2x+6=-4 and -(x+1)=2(3x-1)
Answer by stanbon(75887)

(1/2)x+6 = -4
---
(1/2)x = -10
x = -20
--------------------------

-(x+1) = 2(3x-1)
---
-x-1 = 6x -2
7x = 1
x = 1/7
===============
Cheers,
Stan H.